\Please help me to do part (c)Part a) Draw the shear and bending moment diagrams. Clearly show the values, slopes, and degree of curves (e.g., parabola, 3rPart c) Using the shear and moment diagrams created for the beam shown in “c” determine the maximum bending stress created for the largest positive bending moment and the largest negative bending moment. Consider the beam to be a W 12 x 79.
Answer #1

* Audi 75kn 25kwlo MA Sg Gm B Gm Ay W12429. (4=12in) -0,30487 [I = 662 in I = 0.000275m? & Fy=0) i Ay = 75+ 25 (6) Ay = 225 k22 к. I 150kW От с 6м с бю XSED bmp 6т. 49 . в + 0 т . -450wm . la bota . Poloo - 135o <r - або ем x Bм Р - м 9 то с – 27ooplease rate